SpaceRef

SpaceRef


Chandra Operations CXO Status Report 5 April 2002

Status Report From: Marshall Space Flight Center
Posted: Friday, April 5, 2002

During the last week the loads were interrupted on 2 April due to an execution of the Science Instrument safing SCS 107. The SCS was executed by the OBC as a result of the EPHIN hardware fault flag being set. Analysis of both real-time and dump data showed no problems with the hardware and a procedure was run on 3 April to re-initialize the OBC EPHIN process and resume science. A similar event occurred last December and a software patch to increase two glitch counters was identified to reduce further spurious trips of SCS 107. The patch is expected to be tested and brought for approval for uplink next week. Observations of NGC 6397, BR1202-0725 and HD133880 were impacted and will be rescheduled in an upcoming schedule.

Of note last week was a Chandra image release of the Coma cluster of galaxies. The image shows a concentration of relatively cool gas around two of the larger galaxies that is sustained in a delicate energy balance. For details see: http://chandra.harvard.edu/photo/2002/0150/index.html

The observing schedule for the next week is shown below and includes an observation of GX 3+1 coordinated with RXTE.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.